The minimum value of [ x – 1] + [x – 2] +[x – 3] where     [  ]  represents modulus function is

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: 2
f(x)= |x-1|+|x-2|+|x-3|
f(x)= 6-3x, x< 1
= 4-x, 1\leq x< 2
=x, 2\leq x< 3
=3x-6, x\geq 3
So when you plot the graph of f(x) now, you would see that the minimum value of f(x) lies atx = 2
f(2) = 2
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty